A firm has the following fact pattern for its three product lines.

Product A Product B Product C
Sales Mix 2 1 5
Sales Price Per Unit $84 $40 $22
Variable Cost Per Unit $28 $26 $3

The firm has fixed costs of $1,500,000 per period. How many units of Product C does the firm need to produce to break even?

At the end of your calculations, round up to the nearest whole unit of Product C.
Answer #1

Weighted average contribution margin

= Product A (84-28)*2/8 + Product B (40-26)*1/8 + Product C (22-3)*5/8

= 14 + 1.75 + 11.875

= 27.625

Breakeven point in total = Fixed cost/Weighted average Contribution margin

= 1,500,000/27.625

= 54,299 units

Units of C = 54,299 * 5/8

= 33,937 units